OpenStudy (amyna):

3+2i / 3-2i leave answer in a+bi form the answer is 5/13 + 12/13i , but i keep getting a different answer please help me!!

myininaya (myininaya):

so you have 3+2i/3-2i or do you mean (3+2i)/(3-2i)?

myininaya (myininaya):

i think you mean the latter \[\text{ so we have } \frac{3+2i}{3-2i}\] do you know to put it in a+bi form the first step you must take is to multiply both top and bottom by bottom's conjugate

myininaya (myininaya):

do you know what the conjugate of 3-2i is?

OpenStudy (anonymous):

multipy numerator and denominator with 3+2i

OpenStudy (amyna):

@myininaya yes i multiplied the top and bottom by 3+2i and got 9+6i+6i+4i^2 over 9+6i-6i-4i^2 and got the final answer as 9+12i / 9 but its the wrong answer

myininaya (myininaya):

so how did you get 9+4i^2 is 9?

myininaya (myininaya):

recall i^2 is -1

myininaya (myininaya):

so you got it right so far to say \[\frac{9+6i+6i+4i^2}{9+6i-6i-4i^2}\]

myininaya (myininaya):

and you know again i^2 is -1 so you have \[\frac{9+12i+4(-1)}{9-4(-1)}\] wherever there was an i^2 i just replaced it with -1

myininaya (myininaya):

simplify this and you will see you got the answer as it should be written

OpenStudy (amyna):

Thanks so much @myininaya

myininaya (myininaya):

I think this means you got it so gj